Download
impedance working group update n.
Skip this Video
Loading SlideShow in 5 Seconds..
Impedance working group update PowerPoint Presentation
Download Presentation
Impedance working group update

Impedance working group update

146 Vues Download Presentation
Télécharger la présentation

Impedance working group update

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Impedance working group update Benoit, Carlo, Nicolas for the impedance team 19 Feb 2014

  2. General • Eigenmode simulations with PEC  f, Q and R of the resonator model is enough • Eigenmode simulations with dispersive materials: f, Q and |R| may not be enough as R may now be complex. Request from CST to tell them what to compute and how to compute it. • The broadband resonator frequency has a very strong impact on the beam dynamics  which frequency should be used for LHC? Need to replace this broadband model by actual devices in the near future. • System of equations rewritten for the elliptic/elliptic transitions. • Started simulations of the impedance measurements done in Naples. • Tests done on CERN IT cluster for 1e9 mesh cells: completed in 30 hours  very promising. • Bench measurements: • Current procedure is to measure both with the wire and the probes • Several sets of probes will be built by BE-BI • New improved flanges available with position measurement: very useful!

  3. Machines • PSB • Deadline for ICE meeting on Dec 11th for PSB impedance model and related beam dynamics (very tight, as other priorities took precedence until now)  done • Contribution for TDR (deadline not clear, to be discussed, probably end of year)  done • Coating on ceramic chamber for the new KSW kicker: • resistance is higher than foreseen in the specs (factor 20). • Impedance studied with TLWall, ReWall and CST simulations of the new and present KSW. • However, the new KSW with higher resistance is expected to give an impedance contribution very similar to the present kickers. • A lower resistance will not guarantee that the situation would improve as the transverse impedance peak shifts to lower frequency (from 20 MHz to few MHz range  close to the first possible unstable betatron line). • Ongoing: generate wake model from impedance model (for Headtail) • Ongoing: include numerical form factors for indirect space charge model (from CST simulations). • PS • LIU-PS TDR almost finalized • Recommendation for PS dummy septum released • Localization of impedance seems to point to several locations where wirescanners (64 and 65) and a septum (61) are located. Should be reinvestigated.

  4. Machines • SPS • New model with steps presented at impedance meeting: very good agreement with tune shifts measurements. • Need now to assess the impact on the broadband longitudinal plane • Deadline for draft impedance chapter for the SPS LIU TDR for 28th February (announced on Feb 12th). • Wire scanner measurements completed: idea of Fritz to put an additional wire between the forks seems to reduce some resonances on the wire for some positions of the fork. Simulations ongoing: . • New wire scanner: proposal to install the empty tank to be able to install the scanner system during the run. Simulations show many modes from 666 MHz (very dangerous frequency) of 50 to 120 kOhm. BE-RF/BR is worried and requires more information. • High bandwidth feedback system: request to study the transitions between the 10 stripline kicker modules  almost done, but needs the real length of the transition to conclude. • ZS talk of Hugo at SPSU: measurements on ZS tank. Still many open issues to assess the impedance. Could assess the use of foams to mimic open boundaries.

  5. Machines • LHC • Action on power loss for TCTP collimators and ALFA: bugs seems to be gone in the updated beta version of CST. Tests ongoing with simpler structures. • TDI: LMC decided that cooling will be removed (!), but approved the coating strategy. The hBN blocks will be coated with copper (and NEG), and the stainless steel beam screen will only be coated with NEG. • Crystal collimation: installation approved but not yet the use in MDs. Measurements performed and seem to confirm simulations in parking position (to be presented at the next impedance meeting). Many modes appeared in operating position, indicating that further optimization of the new devices will be necessary. • BGV: cooling may be necessary due to HOMs. Studying air cooling. • TOTEM: measurements performed, and in good agreement with simulations for the longitudinal modes. Presentation at the next impedance meeting. • ALFA: it is likely that the pot was touching the bellow for the measurements of the prototype. Could be the reason why we missed the first mode in measurements.

  6. UA9 Measurements (Joseph, A. Danisi and Olav))

  7. Machines • HL-LHC • All crab cavity modes provided for several options, should implement it in the model. • FCC: • First work on beam screen impedance.

  8. EDMS approvals • Vacuum pilot sector in LSS8 (LHC): need for measurements and guidance for the size of the pickups. • TDI ES and ECR: many comments, in particular on the need to have temperature probes on the jaw. • Second iteration of BGV: comment from Erk JENSEN on 2014-01-28, 08:44: Accepted with reservation (page 15 - impedance: large HOM power to be expected with short bunches and bench measurements must be made and concluded)

  9. Measurements: planning as of 5/02 (from Olav)